### v1

# Mourre theory and spectral analysis of energy-momentum operators in relativistic quantum field theory

Abstract
A central task of theoretical physics is to analyse spectral properties of quantum mechanical observables. In this endeavour, Mourre’s conjugate operator method emerged as an effective tool in the spectral theory of Schrödinger operators. This paper introduces a novel class of examples amenable to Mourre’s method, focusing on the spectra of energy-momentum operators in relativistic quantum field theory. Under the assumption of Lorentz or dilation covariance and the spectrum condition, we provide new proofs of the absolute continuity of the energy-momentum spectra, showcasing the efficiency of Mourre’s method in the relativistic setting.
Keywords: spectral theory, Mourre’s conjugate operator method, absence of singular continuous spectrum, representations of the Poincaré group, dilation-covariant representations
Mathematics Subject Classification 2020: 81Q10
1 Introduction
Understanding spectral properties of quantum mechanical observables is of fundamental importance in theoretical physics. By analysing the spectrum, insights into the system’s stability and long-time evolution can be obtained. Specifically, the spectrum of a self-adjoint operator decomposes into a pure point, absolutely continuous and singular continuous part, where the pure-point part corresponds to bound states and the absolutely continuous part to scattering states. Typically, the singular continuous spectrum, which evades a simple physical interpretation, is empty. The objective of this paper is to prove the absence of singular continuous spectrum for relativistic energy-momentum operators.
Various mathematical techniques have been developed to investigate the spectrum of a self-adjoint operator. One powerful method in this endeavour is Mourre’s conjugate operator method, which is based on a strictly positive commutator estimate. The idea of analysing spectral properties of self-adjoint operators through commutator identities and estimates can be traced back to the pioneering works of Putnam [Pu67], Kato [Ka68], and Lavine [La69]. However, it was Mourre [Mo81] who advanced this approach by introducing local commutator estimates.
Mourre’s method led to substantial progress in the spectral and scattering theory of Schrödinger operators. In his seminal paper, Mourre [Mo81] demonstrated the absence of singular continuous spectrum for 2- and 3-body Schrödinger operators. The applicability of the method to $N$-body systems was extended by Perry, Sigal, and Simon [PSS81] and by Froese and Herbst [FH82]. Mourre’s method also played a decisive role in establishing asymptotic completeness of $N$-body Schrödinger operators (see [DG97] for a textbook exposition). On an abstract level, the mathematical theory underlying the conjugate operator method was notably improved by Amrein, Boutet de Monvel, and Georgescu [ABG96].
While originally developed for non-relativistic quantum mechanics, Mourre’s method has also been extended to other areas. Worthy of note, it has been applied in non-relativistic quantum electrodynamics (QED). Among many works, we mention here that a Mourre estimate was proved by Dereziński and Gérard [DG99] for confined Pauli–Fierz Hamiltonians, by Fröhlich, Griesemer, and Schlein [FGS04] for Compton scattering, by Fröhlich, Griesemer, and Sigal [FGS08] for the standard model of non-realtivistic QED, by Chen, Faupin, Fröhlich, and Sigal [CFFS11] for dressed electrons, and by Møller and Rasmussen [MR13] for the translation-invariant massive Nelson model.
More recently, Mourre’s method has expanded its scope to the domain of Haag–Kastler quantum field theory to prove the existence of asymptotic observables [Kr23]. One of the primary challenges encountered in axiomatic quantum field theory pertains to the abstract nature of the Hamiltonian. Scattering theory enabled the application of Mourre’s method by comparing the abstract dynamics to a more concrete free dynamics.
This paper introduces a novel class of examples amenable to Mourre’s conjugate operator method. We analyse the spectra associated with spacetime translations in quantum field theory, that is, the spectra of the energy-momentum operators $P$. By assuming Lorentz covariance and the spectrum condition, we present a new proof of the well-known theorem asserting that the energy-momentum spectra are purely absolutely continuous if translation-invariant vectors are removed. More precisely, we prove the following proposition, where $\mathcal{H}_{\mathrm{ac}}(A)$ denotes the absolutely continuous spectral subspace of a commuting family of self-adjoint operators $A$ and $x\cdot y=x_{0}y_{0}-\mathbf{x}\cdot\mathbf{y}$, $x,y\in\mathbb{R}^{d}$, denotes the Minkowski scalar product.
Proposition 1.1.
Let $\mathcal{H}$ be a Hilbert space, $U:\mathcal{P}\to\mathfrak{B}(\mathcal{H})$ a strongly continuous unitary representation of the Poincaré group $\mathcal{P}=\mathcal{L}\ltimes\mathbb{R}^{d}$, and $e\in\mathbb{R}^{d}\backslash\{0\}$ a spacetime vector such that $e\cdot e\neq 0$ if $d=2$. Assume that the generators $P=(P_{0},\mathbf{P})$ of the translation subgroup $U|_{\mathbb{R}^{d}}$ obey the spectrum condition (see Assumption 4.1). If $Q_{0}$ denotes the projection onto the subspace of translation-invariant vectors, then
$\mathcal{H}=Q_{0}\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(e\cdot P)=Q_{0}\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(\mathbf{P}).$ (1.1)
The first identity in (1.1) was originally proved by Maison [Ma68, Satz 2] for $d=4$ through the application of Wigner’s theorem. The second identity (absolute continuity of the joint momentum spectrum) has been proved under differing assumptions in the realm of quantum field theory. Buchholz and Fredenhagen [BF82, Proposition 2.2] utilised the locality principle to establish this identity, and Bachmann, Dybalski, and Naaijkens [BDN16, Lemma 4.16] presented a simplified proof under the additional assumption of the existence of a vacuum vector.
We employ Mourre’s method to prove Proposition 1.1, finding it to offer a relatively straightforward proof. As an illustration, we consider the commutation relation between the momentum operator $P_{j}$ and the generator of Lorentz boosts $K_{j}$ in the spatial direction $j$:
$[P_{j},\mathop{}\!\mathrm{i}K_{j}]=P_{0}.$ (1.2)
By virtue of the spectrum condition, this commutator is locally strictly positive on spectral subspaces of $P_{j}$ that are separated from 0. Consequently, the absolute continuity of the spectrum associated with $P_{j}$ emerges as a consequence of the limiting absorption principle (see Theorem 3.5 and Corollary 3.6). The only aspect complicating our proof to a certain extent lies in the regularity requirements of the limiting absorption principle.
An immediate consequence of Proposition 1.1 and the Riemann–Lebesgue lemma is the following clustering property. If $f,g\in\mathcal{H}$ are two arbitrary vectors, then
$\lim_{t\to\pm\infty}\langle f,U(te)g\rangle=\langle f,Q_{0}g\rangle.$ (1.3)
In quantum field theory, the clustering property is well-known for space-like directions (i.e. $e\cdot e<0$) due to the locality principle. It is somewhat unexpected that this property is also valid for light-like (i.e. $e\cdot e=0$) and time-like directions (i.e. $e\cdot e>0$) under the assumption of Lorentz covariance.
In two spacetime dimensions (i.e. $d=2$), the clustering property may not hold for light-like directions. This stems from the potential presence of massless excitations, as elaborated upon in the remark subsequent to Proposition 4.5.
Proposition 1.2.
Let $U$ be as in Proposition 1.1 and let $E_{P_{0}\pm P_{1}}$ denote the spectral measure of $P_{0}\pm P_{1}$. If $d=2$, then
$\mathcal{H}=E_{P_{0}\pm P_{1}}(\{0\})\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(P_{0}\pm P_{1}).$ (1.4)
Another operator of interest in quantum field theory is the mass operator $M=\sqrt{P_{0}^{2}-|\mathbf{P}|^{2}}$. We prove the absence of singular continuous spectrum for the mass operator if $U$ satisfies dilation covariance. A unitary representation $U:\mathbb{R}^{d}\to\mathfrak{B}(\mathcal{H})$ of the translation group $\mathbb{R}^{d}$ is dilation-covariant if a self-adjoint operator $D$ exists such that, for all $t,x\in\mathbb{R}^{d}$,
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tD}U(x)\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tD}=U(\mathop{}\!\mathrm{e}^{-t}x),$ (1.5)
that is, the adjoint action of $\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tD}$ scales the translation $x$ by the factor $\mathop{}\!\mathrm{e}^{-t}$.
Proposition 1.3.
Let $U:\mathbb{R}^{d}\to\mathfrak{B}(\mathcal{H})$ be a strongly continuous unitary representation of the translation group $\mathbb{R}^{d}$. Assume that the generators $P=(P_{0},\mathbf{P})$ of $U$ obey the spectrum condition. If $U$ is dilation-covariant, then the spectrum of the mass operator $M=\sqrt{P_{0}^{2}-|\mathbf{P}|^{2}}$ is purely absolutely continuous in $(0,\infty)$.
Given that the spectrum of the mass operator $M\geq 0$ is contained in the interval $[0,\infty)$, the proposition implies that the pure point spectrum of $M$ must be empty or $\{0\}$, while the singular continuous spectrum is empty. Under the assumptions of Proposition 1.1, the mass operator may have other eigenvalues than 0. The assumption of dilation covariance excludes additional mass shells within the energy-momentum spectrum.
The structure of the paper is as follows. In Section 2, we define locally smooth operators for a commuting family of self-adjoint operators, and we relate the existence of a locally smooth operator to the regularity of the joint spectrum. In Section 3, we review essential results of Mourre’s conjugate operator method. We explain how to obtain locally smooth operators based on a strictly positive commutator estimate. In Section 4, we present the proofs of Propositions 1.11.3.
2 Locally smooth operators
Let $U:\mathbb{R}^{n}\to\mathfrak{B}(\mathcal{H})$ be a strongly continuous unitary representation of $\mathbb{R}^{n}$ on a Hilbert space $\mathcal{H}$ with self-adjoint generators $H=(H_{1},\dots,H_{n})$. We denote by $D(H)$ the intersection of the domains of $H_{1},\dots,H_{n}$, by $E$ the spectral measure of the family $H$, and by $\sigma(H)\subset\mathbb{R}^{n}$ its joint spectrum. We consider $D(H)$ as a Banach space equipped with the graph norm.
Definition 2.1.
A continuous operator $T:D(H)\to\mathcal{H}$ is locally $H$-smooth on an open set $J\subset\mathbb{R}^{n}$ if, for every compact subset $K\subset J$, a constant $C_{K}$ exists such that, for all $f\in\mathcal{H}$,
$\int_{\mathbb{R}^{n}}\|TU(x)E(K)f\|^{2}\mathop{}\!\mathrm{d}x\leq C_{K}\|f\|^{2}.$ (2.1)
In [Kr23, Definition A.9], we defined locally $H$-smooth operators by demanding that all subfamilies of $H$ are locally $H$-smooth in the sense of the above definition. In this paper, we prefer the weaker formulation of Definition 2.1 because it is sufficient to characterise the regularity of the joint spectrum $\sigma(H)$. The following proposition is a straightforward generalisation of [RS78, Theorem XIII.23].
Proposition 2.2.
If $T$ is locally $H$-smooth on $J$, then $E(J)\overline{\mathop{}\!\mathrm{ran}(T^{*})}\subset E(J)\mathcal{H}_{\mathrm{ac}}(H)$. If, additionally, $\ker(T)=\{0\}$, then the joint spectrum of $H$ is purely absolutely continuous in $J$.
Proof.
Let $f\in E(J)\overline{\mathop{}\!\mathrm{ran}(T^{*})}$. Due to the inner regularity of the spectral measure $E$, we can approximate $f$ by a sequence $\{f_{j}\}_{j\in\mathbb{N}}$ such that, for every $j\in\mathbb{N}$, $f_{j}\in E(K_{j})\mathop{}\!\mathrm{ran}(T^{*})$ for a compact subset $K_{j}\subset J$. If, for every $j\in\mathbb{N}$, $f_{j}\in E(J)\mathcal{H}_{\mathrm{ac}}(H)$, then also $f\in E(J)\mathcal{H}_{\mathrm{ac}}(H)$ because $E(J)\mathcal{H}_{\mathrm{ac}}(H)$ is a closed subspace of $\mathcal{H}$. We write $f_{j}=E(K_{j})T^{*}g_{j}$ with $g_{j}\in D(T^{*})$. Let $\mu_{f_{j}}$ be the spectral measure associated with $f_{j}$ and $F_{j}$ its inverse Fourier transform:
$F_{j}(x)=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}p\cdot x}\mathop{}\!\mathrm{d}\mu_{f_{j}}(p)=\frac{1}{(2\pi)^{n}}\langle E(K_{j})T^{*}g_{j},U(x)f_{j}\rangle.$ (2.2)
The function $F_{j}$ is square-integrable because $|F_{j}(x)|\leq(2\pi)^{-n}\|g_{j}\|\|TU(x)E(K_{j})f_{j}\|$ and $T$ is locally $H$-smooth on $J$ by assumption. Thus, $\mathop{}\!\mathrm{d}\mu_{f_{j}}(p)=\hat{F}_{j}(p)\mathop{}\!\mathrm{d}p$ is an absolutely continuous measure. To prove the second statement, we apply the decomposition $\mathcal{H}=\ker(T)\oplus\overline{\mathop{}\!\mathrm{ran}(T^{*})}$. If $\ker(T)=\{0\}$, then $E(J)\mathcal{H}=E(J)\overline{\mathop{}\!\mathrm{ran}(T^{*})}\subset E(J)\mathcal{H}_{\mathrm{ac}}(H)$. We conclude $E(J)\mathcal{H}=E(J)\mathcal{H}_{\mathrm{ac}}(H)$ because $\mathcal{H}_{\mathrm{ac}}(H)\subset\mathcal{H}$, that is, the joint spectrum of $H$ is purely absolutely continuous in $J$. ∎
It is possible to determine the optimal constant $C_{K}$ for the bound (2.1), which is useful to verify the $H$-smoothness of an operator $T$. The following proposition is proved as in [Kr23, Proposition A.10].
Proposition 2.3.
If $T$ is locally $H$-smooth, then the optimal constant for the bound (2.1) is
$C_{K}^{0}=2^{n}\sup_{\lambda\in\mathbb{R}^{n},\mu\in(0,1)^{n}}\mu_{1}\dots\mu_{n}\|TE(K)R(\lambda+\mathop{}\!\mathrm{i}\mu)\|^{2},$ (2.3)
where
$R(\lambda+\mathop{}\!\mathrm{i}\mu)=\prod_{j=1}^{n}(H_{j}-\lambda_{j}-\mathop{}\!\mathrm{i}\mu_{j})^{-1}.$ (2.4)
Moreover, a continuous operator $T:D(H)\to\mathcal{H}$ is locally $H$-smooth on $J$ if $C_{K}^{0}<\infty$ for every compact subset $K\subset J$.
If $n=1$, we can formulate a necessary and sufficient condition for $C_{K}^{0}<\infty$ in terms of a limiting absorption principle (see Section 3).
Proposition 2.4.
Let $n=1$. For every compact subset $K\subset\mathbb{R}$, $C_{K}^{0}<\infty$ if and only if
$\sup_{\lambda\in K,\mu\in(0,1)}\|T\Im R(\lambda+\mathop{}\!\mathrm{i}\mu)T^{*}\|<\infty.$ (2.5)
The proof of Proposition 2.4 is contained in the proof of [ABG96, Proposition 7.1.1]; see [Kr23, Theorem A.12] for a generalisation to $n\geq 2$. In the next section, we apply Proposition 2.4 to obtain locally smooth operators.
3 Mourre’s conjugate operator method
Mourre’s conjugate operator method is a mathematical tool to analyse the spectrum of a self-adjoint operator based on a commutator estimate. In this section, we outline the framework and state the most important results of the method, in particular the limiting absorption principle. At the end of this section, we apply Mourre’s method to construct locally smooth operators.
We introduce the following regularity classes, which are relevant for defining commutators.
Definition 3.1.
Let $A$ be a self-adjoint operator on $\mathcal{H}$ and let $k\in\mathbb{N}\cup\{\infty\}$. We denote by $C^{k}(A)$ the space of all self-adjoint operators $H$ such that $t\mapsto\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tA}(H+\mathop{}\!\mathrm{i})^{-1}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tA}$ is a $C^{k}$-map in the strong operator topology. We denote by $C_{\mathrm{u}}^{k}(A)$ the subspace of operators $H$ for which the same map is $C^{k}$ in norm.
We collect some properties of the class $C^{1}(A)$ from [ABG96, Theorem 6.2.10]. If $H\in C^{1}(A)$, then $D(A)\cap D(H)$ is dense in $D(H)$ (equipped with the graph topology). Moreover, the sesquilinear form defined by the commutator $HA-AH$ on $D(A)\cap D(H)$ extends to a continuous sesquilinear form on $D(H)$. We denote the extended sesquilinear form by $[H,A]$. It is possible to characterise the class $C^{1}(A)$ in terms of the commutator $[H,A]$.
Proposition 3.2 ([ABG96, Theorem 6.2.10 (a)]).
A self-adjoint operator $H$ belongs to the class $C^{1}(A)$ if and only if there exists $z\in\rho(H)$ such that $\{f\in D(A)\mid(H-z)^{-1}f\in D(A)\text{ and }(H-\overline{z})^{-1}f\in D(A)\}$ is a core for $A$ and, for all $f\in D(A)\cap D(H)$,
$|\langle Hf,Af\rangle-\langle Af,Hf\rangle|\leq c(\|Hf\|^{2}+\|f\|^{2}).$ (3.1)
Having formalised the commutator of two self-adjoint unbounded operators, we are able to define strictly positive commutator estimates.
Definition 3.3.
A self-adjoint operator $H$ obeys a Mourre estimate on an open and bounded subset $J\subset\mathbb{R}$ if a self-adjoint operator $A$ (conjugate operator) exists such that $H\in C^{1}(A)$ and, for an $a>0$,
$E(J)[H,\mathop{}\!\mathrm{i}A]E(J)\geq aE(J),$ (3.2)
where $E$ is the spectral measure of $H$.
Example.
Let $H$ be $A$-homogeneous, that is, for every $t,x\in\mathbb{R}$,
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tA}\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}xH}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tA}=\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}\mathop{}\!\mathrm{e}^{-t}xH}.$ (3.3)
It is not difficult to show that (3.3) implies $H\in C^{\infty}(A)$ and $[H,\mathop{}\!\mathrm{i}A]=H$. Moreover, for every open and bounded interval $J=(a,b)\subset(0,\infty)$, $0<a<b<\infty$,
$E(J)[H,\mathop{}\!\mathrm{i}A]E(J)=HE(J)\geq aE(J),$ (3.4)
that is, $H$ obeys a Mourre estimate on $J$ with conjugate operator $A$. Similarly, for every open and bounded interval $I=(a,b)\subset(-\infty,0)$, $-\infty<a<b<0$,
$E(I)[H,-\mathop{}\!\mathrm{i}A]E(I)=-HE(I)\geq-bE(I),$ (3.5)
that is, $H$ obeys a Mourre estimate on $I$ with conjugate operator $-A$.
A consequence of the Mourre estimate on $J$ is that $H$ has no eigenvalues in $J$. This fact emerges as a direct corollary of the virial theorem [ABG96, Proposition 7.2.10].
Theorem 3.4 (Virial theorem).
Let $H\in C^{1}(A)$. If $f$ is an eigenvector of $H$, then $\langle f,[H,A]f\rangle=0$.
Next, we state the limiting absorption principle. To formulate the limiting absorption principle under optimal conditions, we introduce new regularity classes. Let $p\in[1,\infty]$ and $k\in\mathbb{N}$. For $0<s<k$, we denote by $\mathcal{C}^{s,p}(A)$ the real interpolation space [ABG96, (5.2.6)]
$\mathcal{C}^{s,p}(A)=(C^{k}_{\mathrm{u}}(A),C^{0}_{\mathrm{u}}(A))_{\theta,p},\ \theta=1-s/k.$ (3.6)
Moreover, we write $H\in C^{k+0}(A)$ [ABG96, p. 204] if $H\in C^{k}(A)$ and if the operator-valued function $S(t)=\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tA}\mathrm{ad}_{A}^{k}((H+\mathop{}\!\mathrm{i})^{-1})\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tA}$ is Dini continuous in norm, that is,
$\int_{|t|<1}\|S(t)-S(0)\|\frac{\mathop{}\!\mathrm{d}t}{t}<\infty.$ (3.7)
We mention the following chain of inclusions [ABG96, (5.2.14)]:
$C^{k+1}(A)\subset C^{k+0}(A)\subset\mathcal{C}^{k,1}(A)\subset C^{k}_{\mathrm{u}}(A),\ k\in\mathbb{N}.$ (3.8)
Note that $C^{k+1}(A)\subset C^{k+0}(A)$ as a consequence of the mean value theorem and the uniform boundedness principle.
We say that $H$ has a spectral gap if $\sigma(H)\neq\mathbb{R}$ and we write $\langle A\rangle$ for $\sqrt{A^{2}+1}$.
Theorem 3.5 (Limiting absorption principle).
Let $H\in\mathcal{C}^{1,1}(A)$ if $H$ has a spectral gap and $H\in C^{1+0}(A)$ if $\sigma(H)=\mathbb{R}$. If $H$ obeys a Mourre estimate on $J$, then, for every compact subset $K\subset J$ and every $\nu>1/2$,
$\sup_{\lambda\in K,\mu>0}\|\langle A\rangle^{-\nu}(H-\lambda\mp\mathop{}\!\mathrm{i}\mu)^{-1}\langle A\rangle^{-\nu}\|<\infty.$ (3.9)
Remark.
The assumption $H\in\mathcal{C}^{1,1}(A)$ is optimal on the Besov scale in the sense that counterexamples of the limiting absorption principle for operators possessing less regularity exist [ABG96, Section 7.B]. A proof of the limiting absorption principle under the assumption $H\in\mathcal{C}^{1,1}(A)$ has been accomplished in the case that $H$ has a spectral gap [ABG96, Theorem 7.4.1]. Sahbani [Sa97] proved the limiting absorption principle under the more restrictive assumption $H\in C^{1+0}(A)$. It is an open problem to prove the limiting absorption principle for $H\in\mathcal{C}^{1,1}(A)$ if $\sigma(H)=\mathbb{R}$.
According to Proposition 2.2, $H$ has no singular spectrum in $J$ if an injective locally $H$-smooth operator on $J$ exists. If each $H_{i}$ obeys a Mourre estimate with conjugate operator $A_{i}$ such that, for $i\neq j$, $A_{i}$ commutes strongly with $H_{j}$, then, for every $\nu>1/2$, $\langle A_{1}\rangle^{-\nu}\dots\langle A_{n}\rangle^{-\nu}$ is an injective operator that is locally $H$-smooth. This is a corollary of the limiting absorption principle and the results from Section 2.
Corollary 3.6.
Consider a commuting family of self-adjoint operators $H=(H_{1},\dots,H_{n})$ and a corresponding family $A=(A_{1},\dots,A_{n})$ of self-adjoint operators such that, for every $i\in\{1,\dots,n\}$, $H_{i}\in\mathcal{C}^{1,1}(A_{i})$ if $H_{i}$ has a spectral gap and $H_{i}\in C^{1+0}(A_{i})$ if $\sigma(H)=\mathbb{R}$. If $H_{i}$ obeys a Mourre estimate on $J_{i}$ with conjugate operator $A_{i}$ and, for $i\neq j$, $A_{i}$ commutes strongly with $H_{j}$, then the joint spectrum of $H$ is purely absolutely continuous in the product region $J=J_{1}\times\dots\times J_{n}$.
Proof.
We demonstrate that the injective operator $T=\langle A_{1}\rangle^{-\nu}\dots\langle A_{n}\rangle^{-\nu}$ is locally $H$-smooth on $J$ for every $\nu>1/2$. If this claim is proved, the statement follows from Proposition 2.2. We show that the optimal constant $C_{K}^{0}$ from Proposition 2.3 is finite for every compact subset $K\subset J$. It suffices to consider compact hyperrectangles $K=K_{1}\times\dots\times K_{n}$, where $K_{i}\subset J_{i}$ are compact intervals because every compact subset $K\subset J$ can be covered by finitely many compact hyperrectangles in $J$. If $E_{i}$ denotes the spectral measure of $H_{i}$, then
$E(K)=E_{1}(K_{1})\dots E_{n}(K_{n}).$ (3.10)
By assumption, $A_{i}$ commutes strongly with $H_{j}$ if $i\neq j$; hence,
$C_{K}^{0}$ $=2^{n}\sup_{\lambda\in\mathbb{R}^{n},\mu\in(0,1)^{n}}\mu_{1}\dots\mu_{n}\|\langle A_{1}\rangle^{-\nu}\dots\langle A_{n}\rangle^{-\nu}E(K)R(\lambda+\mathop{}\!\mathrm{i}\mu)\|^{2}$
$\leq\prod_{i=1}^{n}2\sup_{\lambda_{i}\in\mathbb{R},\mu_{i}\in(0,1)}\mu_{i}\|\langle A_{i}\rangle^{-\nu}E_{i}(K_{i})R_{i}(\lambda_{i}+\mathop{}\!\mathrm{i}\mu_{i})\|^{2}.$ (3.11)
Each factor on the r.h.s. is finite due to Proposition 2.4 and the limiting absorption principle (Theorem 3.5). ∎
4 Spectral analysis of relativistic energy-momentum operators
We analyse the spectra associated with the generators of spacetime translations, employing the methods presented above. In the first subsection, we analyse unitary representations of the Poincaré group. In the second subsection, we explore dilation-covariant representations of the translation group.
4.1 Representations of the Poincaré group
The Poincaré group is the semi-direct product $\mathcal{P}=\mathcal{L}\ltimes\mathbb{R}^{d}$ of the Lorentz group $\mathcal{L}=O(d-1,1)$ and the translation group $\mathbb{R}^{d}$. Its multiplication law is defined as follows:
$(\lambda_{1},a_{1})\cdot(\lambda_{2},a_{2})=(\lambda_{1}\lambda_{2},a_{1}+\lambda_{1}a_{2}),\ \lambda_{1},\lambda_{2}\in\mathcal{L},a_{1},a_{2}\in\mathbb{R}^{d}.$ (4.1)
Let $U:\mathcal{P}\to\mathfrak{B}(\mathcal{H})$ be a strongly continuous unitary representation of the Poincaré group on a Hilbert space $\mathcal{H}$. The generators of the translation subgroup $U|_{\mathbb{R}^{d}}$ are the energy-momentum operators $P=(P_{0},\mathbf{P})$. We denote by $\Lambda_{1}(t),\dots,\Lambda_{s}(t)$ the Lorentz boosts in the space directions $11,\dots,s=d-$ 1, for example,
$\Lambda_{1}(t)=\begin{pmatrix}\cosh(t)&\sinh(t)&0&\cdots&0\\ \sinh(t)&\cosh(t)&0&\cdots&0\\ 0&0&1&&0\\ \vdots&\vdots&&\ddots&\\ 0&0&0&&1\end{pmatrix},$ (4.7)
and by $K_{1},\dots,K_{s}$ the self-adjoint generators of the Lorentz boosts, that is, $U(\Lambda_{j}(t),0)=\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{j}}$, $j\in\{1,\dots,s\}$. We mention the following identities, which are consequences of the multiplication law ($t,x\in\mathbb{R}$):
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{j}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}xP_{0}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tK_{j}}$ $=\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}x(\cosh(t)P_{0}-\sinh(t)P_{j})},$ (4.8)
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{j}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}xP_{j}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tK_{j}}$ $=\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}x(\cosh(t)P_{j}-\sinh(t)P_{0})}.$ (4.9)
Formally, by differentiating in $x=0$ and $t=0$, these identities are equivalent to the commutation relations $[P_{0},\mathop{}\!\mathrm{i}K_{j}]=P_{j}$ and $[P_{j},\mathop{}\!\mathrm{i}K_{j}]=P_{0}$.
Utilising Wigner’s theorem, it has been proved in [Ma68, Satz 2] that the spectra of the energy-momentum operators are absolutely continuous if translation-invariant vectors are removed. Translation-invariant vectors are vectors $f\in\mathcal{H}$ for which $U(1,x)f=f$ for all $x\in\mathbb{R}^{d}$. We present a new proof employing Mourre’s conjugate operator method. Our analysis crucially depends on the spectrum condition, which will be assumed for the remainder of this section.
Assumption 4.1 (Spectrum condition).
The joint spectrum of the energy-momentum operators is a subset of the closed forward light cone $V_{+}=\{p=(p_{0},\mathbf{p})\in\mathbb{R}^{d}\mid p_{0}\geq|\mathbf{p}|\}$ (i.e. $\sigma(P)\subset V_{+}$).
The spectrum condition is a quantum field theory axiom in the frameworks of Wightman and Haag–Kastler. From the spectrum condition, it follows that the momentum operators are relatively bounded w.r.t. the energy operator $P_{0}$, that is, $|\mathbf{P}|\leq P_{0}$.
$p_{0}$$\mathbf{p}$$S_{\lambda}$$\lambda$
Figure 1: The set $S_{\lambda}\subset V_{+}$ is invariant under Lorentz boosts.
We introduce, for $\lambda>0$, the following Lorentz-invariant sets (see Figure 1):
$S_{\lambda}=\{\Lambda_{1}(t_{1})\dots\Lambda_{s}(t_{s})(p_{0},\mathbf{0})\mid t_{1},\dots,t_{s}\in\mathbb{R},p_{0}\in[0,\lambda]\}.$ (4.10)
Clearly, the sets $S_{\lambda}$ cover the forward light cone $V_{+}$ (i.e. $\bigcup_{\lambda>0}S_{\lambda}=V_{+}$). Hence, by the spectrum condition, the subspaces $E(S_{\lambda})\mathcal{H}$ cover the Hilbert space $\mathcal{H}$, where $E$ is the joint spectral measure of the energy-momentum operators $P$. Moreover,
$\mathcal{H}_{\mathrm{ac}}(\mathbf{P})=\bigcup_{\lambda>0}\mathcal{H}_{\mathrm{ac}}(\mathbf{P}_{\lambda}),\ \mathcal{H}_{\mathrm{ac}}(P_{0})=\bigcup_{\lambda>0}\mathcal{H}_{\mathrm{ac}}(P_{0,\lambda}),$ (4.11)
where $\mathbf{P}_{\lambda}=E(S_{\lambda})\mathbf{P}$ and $P_{0,\lambda}=E(S_{\lambda})P_{0}$ are operators on $E(S_{\lambda})\mathcal{H}$. The advantage of introducing the operators $P_{0,\lambda}$ and $\mathbf{P}_{\lambda}$ is that $P_{0,\lambda}$ is relatively bounded w.r.t. $P_{j,\lambda}=E(S_{\lambda})P_{j}$.
The following lemmas will be applied to prove absolute continuity of the energy-momentum spectra.
Lemma 4.2.
For every $\lambda>0$ and $j\in\{1,\dots,s\}$, $P_{0,\lambda}\in C^{\infty}(K_{j})$, $P_{j,\lambda}\in C^{\infty}(K_{j})$, and
$[P_{0,\lambda},\mathop{}\!\mathrm{i}K_{j}]=P_{j,\lambda},$ (4.12)
$[P_{j,\lambda},\mathop{}\!\mathrm{i}K_{j}]=P_{0,\lambda}.$ (4.13)
Proof.
The spectral projection $E(S_{\lambda})$ commutes with $K_{j}$ because $S_{\lambda}$ is a Lorentz-invariant set. Thus, we can replace $P_{0}$ and $P_{j}$ in (4.9) with $P_{0,\lambda}$ and $P_{j,\lambda}$:
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{j}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}xP_{j,\lambda}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tK_{j}}$ $=\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}x(\cosh(t)P_{j,\lambda}-\sinh(t)P_{0,\lambda})}.$ (4.14)
From (4.14), we obtain the following identity:
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{j}}(P_{j,\lambda}+\mathop{}\!\mathrm{i})^{-1}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tK_{j}}$ $=(\cosh(t)P_{j,\lambda}-\sinh(t)P_{0,\lambda}+\mathop{}\!\mathrm{i})^{-1}.$ (4.15)
The r.h.s. is a smooth function in $t$ in the strong operator topology because $P_{0,\lambda}$ and $P_{j,\lambda}$ are relatively bounded w.r.t. each other; hence, $P_{j,\lambda}\in C^{\infty}(K_{j})$. Moreover, by differentiating (4.14) in $t=0$ and $x=0$, we obtain, for every $f,g\in D(P_{j,\lambda})\cap D(K_{j})$,
$\langle P_{j,\lambda}f,K_{j}g\rangle-\langle K_{j}f,P_{j,\lambda}g\rangle=-\mathop{}\!\mathrm{i}\langle f,P_{0,\lambda}g\rangle.$ (4.16)
Because $P_{j,\lambda}\in C^{1}(K_{j})$, the sesquilinear form (4.16) has a unique extension to $D(P_{j,\lambda})=D(P_{0,\lambda})$, yielding the commutator identity $[P_{j,\lambda},\mathop{}\!\mathrm{i}K_{j}]=P_{0,\lambda}$. The proof of $P_{0,\lambda}\in C^{\infty}(K_{j})$ and $[P_{0,\lambda},\mathop{}\!\mathrm{i}K_{j}]=P_{j,\lambda}$ is analogous. ∎
Lemma 4.3.
Let $f\in\mathcal{H}$. If $P_{\mu}f=0$ for one $\mu\in\{0,\dots,s\}$, then $P_{\mu}f=0$ for all $\mu\in\{0,\dots,s\}$, that is, if a vector is translation-invariant in one spacetime direction, then it is translation-invariant in all spacetime directions.
Proof.
If $P_{0}f=0$, then $P_{j}f=0$ for all $j\in\{1,\dots,s\}$ due to the spectrum condition. If $P_{j}f=0$ for one $j\in\{1,\dots,s\}$, then also $P_{j,\lambda}f=0$ for all $\lambda>0$. From the commutation relation $[P_{j,\lambda},\mathop{}\!\mathrm{i}K_{j}]=P_{0,\lambda}$ and the virial theorem (Theorem 3.4), it follows that
$\|\sqrt{P_{0,\lambda}}f\|^{2}=\langle f,[P_{j,\lambda},\mathop{}\!\mathrm{i}K_{j}]f\rangle=0;$ (4.17)
hence, $P_{0,\lambda}f=0$ for all $\lambda>0$, that is, $P_{0}f=0$. ∎
We split the statements of Proposition 1.1 and Proposition 1.2 into three separate propositions. First, we prove the absolute continuity of the joint momentum spectrum, which implies also the absolute continuity of $e\cdot P$ for space-like vectors $e\in\mathbb{R}^{d}$. Subsequently, we prove the absolute continuity of $e\cdot P$ for light-like and time-like vectors $e$, respectively. In the latter cases, $e\cdot P$ is unitarily equivalent to a scalar multiple of $P_{0}-P_{1}$, $P_{0}+P_{1}$, or $P_{0}$, depending on whether $e$ is light-like or time-like. We denote by $Q_{0}$ the projection onto the subspace of translation-invariant vectors.
Proposition 4.4.
Under the assumptions of Proposition 1.1, $\mathcal{H}=Q_{0}\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(\mathbf{P})$.
Proof.
Let $a>0$ and let $J$ be an open and bounded subset of $(-\infty,-a]\cup[a,\infty)$. From Lemma 4.2 and the spectrum condition, we obtain the following Mourre estimate on the Hilbert space $E(S_{\lambda})\mathcal{H}$:
$E_{j,\lambda}(J)[P_{j,\lambda},\mathop{}\!\mathrm{i}K_{j}]E_{j,\lambda}(J)=P_{0,\lambda}E_{j,\lambda}(J)\geq aE_{j,\lambda}(J),$ (4.18)
where $E_{j,\lambda}$ denotes the spectral measure of $P_{j,\lambda}$. Because $a>0$ can be arbitrarily small, it follows from Corollary 3.6 that $E(\{p\in S_{\lambda}\mid\forall j\in\{1,\dots,s\}\colon p_{j}\neq 0\})\mathcal{H}\subset\mathcal{H}_{\mathrm{ac}}(\mathbf{P}_{\lambda})$. To apply Corollary 3.6, we used that $P_{j,\lambda}\in C^{\infty}(K_{j})\subset C^{1+0}(K_{j})$ due to (3.8). By taking the union over $\lambda>0$, we obtain $E(\{p\in V_{+}\mid\forall j\in\{1,\dots,s\}\colon p_{j}\neq 0\})\mathcal{H}\subset\mathcal{H}_{\mathrm{ac}}(\mathbf{P})$. By Lemma 4.3, $E(\{p\in V_{+}\mid\forall j\in\{1,\dots,s\}\colon p_{j}\neq 0\})\mathcal{H}=E(V_{+}\backslash\{0\})\mathcal{H}$. The statement follows from the decomposition $\mathcal{H}=Q_{0}\mathcal{H}\oplus E(V_{+}\backslash\{0\})\mathcal{H}$. ∎
Proposition 4.5.
Under the assumptions of Proposition 1.1, $\mathcal{H}=Q_{0}\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(P_{0}\pm P_{1})$ if $d\geq 3$ and $\mathcal{H}=E_{P_{0}\pm P_{1}}(\{0\})\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(P_{0}\pm P_{1})$ if $d=2$.
Proof.
From (4.8) and (4.9), we obtain the following identity:
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{1}}\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}x(P_{0}\pm P_{1})}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tK_{1}}=\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}\mathop{}\!\mathrm{e}^{\mp t}x(P_{0}\pm P_{1})}.$ (4.19)
In the terminology of the example subsequent to Definition 3.3, the operator $P_{0}\pm P_{1}$ is ±$\pm K_{1}$-homogeneous. It follows that $P_{0}\pm P_{1}$ obeys a Mourre estimate on every open and bounded interval that is separated from 0. Thus, by Corollary 3.6, $E(V_{+}\backslash\{p_{0}\pm p_{1}\neq 0\})\mathcal{H}\subset\mathcal{H}_{\mathrm{ac}}(P_{0}\pm P_{1})$. If $d\geq 3$, then $E(V_{+}\backslash\{p_{0}\pm p_{1}\neq 0\})\mathcal{H}=E(V_{+}\backslash\{0\})\mathcal{H}$. In fact, for $f\in\mathcal{H}$, $P_{0}f=\pm P_{1}f$ implies $P_{2}f=0$ by the spectrum condition. And, by Lemma 4.3, $P_{2}f=0$ implies $P_{0}f=P_{1}f=0$. ∎
Remark.
Assume that an eigenstate $f\in\mathcal{H}/Q_{0}\mathcal{H}$ with eigenvalue 0 of the mass operator $M=\sqrt{P_{0}^{2}-|\mathbf{P}|^{2}}$ exists. If $d=2$, then $P_{0}f=|\mathbf{P}|f=|P_{1}|f$ for such a massless excitation. If $f_{\pm}$ denotes the positive/negative momentum component of $f$, then $(P_{0}\mp P_{1})f_{\pm}=0$. This illustrates that, in the case $d=2$, the subspace $E_{P_{0}\pm P_{1}}(\{0\})\mathcal{H}$ may be larger than $Q_{0}\mathcal{H}$.
Proposition 4.6.
Under the assumptions of Proposition 1.1, $\mathcal{H}=Q_{0}\mathcal{H}\oplus\mathcal{H}_{\mathrm{ac}}(P_{0})$.
Proof.
Let $\lambda>0$, select $\theta\in C_{c}^{\infty}(\lambda,\infty)$, and set $F(P_{\lambda})=\theta(P_{0,\lambda})/P_{1,\lambda}$. The operator $F(P_{\lambda})$ is well-defined because if $p\in S_{\lambda}$ with $p_{0}\in\mathrm{supp}(\theta)$, then $p_{1}$ is separated from 0. We define the following operator on the domain $D(A_{\lambda})=D(K_{1})$:
$A_{\lambda}=\frac{1}{2}(F(P_{\lambda})K_{1}+K_{1}F(P_{\lambda})).$ (4.20)
By Lemma 4.7 below, the operator $A_{\lambda}$ is essentially self-adjoint, $P_{0,\lambda}\in C^{\infty}(\overline{A_{\lambda}})$, and, for $0<a<b$,
$E_{0,\lambda}(\lambda+(a,b))[P_{0,\lambda},\mathop{}\!\mathrm{i}\overline{A_{\lambda}}]E_{0,\lambda}(\lambda+(a,b))=\theta(P_{0,\lambda})E_{0,\lambda}(\lambda+(a,b)),$ (4.21)
where $E_{0,\lambda}$ is the spectral measure of $P_{0,\lambda}$. If we select $\theta\in C_{c}^{\infty}(\lambda,\infty)$ such that $\theta=1$ on $\lambda+(a,b)$, then $P_{0,\lambda}$ obeys a Mourre estimate on $\lambda+(a,b)$. From Corollary 3.6, it follows that $E(\{p\in S_{\lambda}\mid p_{0}\in\lambda+(a,b)\})\mathcal{H}\subset\mathcal{H}_{\mathrm{ac}}(P_{0,\lambda})$. We take the union over $\lambda$, $a$, and $b$:
$E(V_{+}\backslash\{\mathbf{p}=0\})\mathcal{H}=\bigcup_{\lambda>0,b>a>0}E(\{p\in S_{\lambda}\mid p_{0}\in\lambda+(a,b)\})\mathcal{H}\subset\mathcal{H}_{\mathrm{ac}}(P_{0}).$ (4.22)
By Lemma 4.3, the l.h.s. is equal to $E(V_{+}\backslash\{0\})\mathcal{H}$. ∎
Remark.
The choice of the conjugate operator $\overline{A_{\lambda}}$ is similar as in [ABG96, Lemma 7.6.4].
Lemma 4.7.
Let $\lambda>0$, $\theta\in C_{c}^{\infty}(\lambda,\infty)$, and set $F(P_{\lambda})=\theta(P_{0,\lambda})/P_{1,\lambda}$. The symmetric operator
$A_{\lambda}=\frac{1}{2}(F(P_{\lambda})K_{1}+K_{1}F(P_{\lambda}))$ (4.23)
is essentially self-adjoint on $D(K_{1})$. If we denote its self-adjoint closure by $\overline{A_{\lambda}}$, then $P_{0,\lambda}\in C^{\infty}(\overline{A_{\lambda}})$ and
$[P_{0,\lambda},\mathop{}\!\mathrm{i}\overline{A_{\lambda}}]=\theta(P_{0,\lambda}).$ (4.24)
Proof.
  1. (i)
    If $G\in C_{c}^{\infty}(\mathbb{R}^{d},\mathbb{R})$ and $j\in\{1,\dots,s\}$, then $G(P_{\lambda})\in C^{\infty}(K_{j})$ and the following commutator identity holds:
    $[G(P_{\lambda}),\mathop{}\!\mathrm{i}K_{j}]$ $=-\partial_{t}\big{|}_{0}\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{j}}G(P_{\lambda})\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tK_{j}}$
    $=-\partial_{t}\big{|}_{0}G(\cosh(t)P_{0,\lambda}-\sinh(t)P_{j,\lambda},P_{1,\lambda},\dots,\cosh(t)P_{j,\lambda}-\sinh(t)P_{0,\lambda},\dots,P_{s,\lambda})$
    $=\partial_{0}G(P_{\lambda})P_{j,\lambda}+\partial_{j}G(P_{\lambda})P_{0,\lambda}.$ (4.25)
    Applying this identity to $F(P_{\lambda})$, we obtain the following commutator:
    $[F(P_{\lambda}),\mathop{}\!\mathrm{i}K_{1}]=\theta^{\prime}(P_{0,\lambda})-\theta(P_{0,\lambda})\frac{P_{0,\lambda}}{P_{1,\lambda}^{2}}.$ (4.26)
    The operator on the r.h.s. is bounded. It follows that $F(P_{\lambda})$ leaves $D(K_{1})$ invariant. In particular, $A_{\lambda}$ is well-defined on $D(K_{1})$.
  2. (ii)
    We apply Nelson’s commutator theorem [RS75, Theorem X.36] to establish the essential self-adjointness of $A_{\lambda}$ on $D(K_{1})$. Setting $N=K_{1}^{2}+1$, we define the integer scale $\mathcal{H}_{k}$, $k\in\mathbb{Z}$, corresponding to $N$ as the completion of $D(N^{k/2})$ with respect to the norm $\|f\|_{k}=\|N^{k/2}f\|$. Clearly, $A_{\lambda}$ is a symmetric bounded operator from $\mathcal{H}_{n}$ to $\mathcal{H}_{-n}$ for every $n\geq 1$. Moreover, as an operator identity from $\mathcal{H}_{n+2}$ to $\mathcal{H}_{-n-2}$,
    $NA_{\lambda}-A_{\lambda}N=2K_{1}[K_{1},F(P_{\lambda})]K_{1}+\frac{1}{2}[K_{1},[K_{1},[K_{1},F(P_{\lambda})]]].$ (4.27)
    The triple commutator in the second summand is a bounded operator on $\mathcal{H}$ according to (4.25). It follows that, for every $f\in\mathcal{H}_{n+2}$,
    $\|NA_{\lambda}f-A_{\lambda}Nf\|_{-n}\leq c\|f\|_{n};$ (4.28)
    hence, the commutator $NA_{\lambda}-A_{\lambda}N$ extends to a bounded operator from $\mathcal{H}_{n}$ to $\mathcal{H}_{-n}$. We conclude that $A_{\lambda}$ is essentially self-adjoint on any core of $N$, particularly on $D(K_{1}^{2})$. The closure of $A_{\lambda}$ restricted to $D(K_{1}^{2})$ coincides with the closure of $A_{\lambda}$ defined on $D(K_{1})$. Thus, $A_{\lambda}$ is essentially self-adjoint on $D(K_{1})$.
  3. (iii)
    If $\chi\in C_{c}^{\infty}(\mathbb{R})$ is a function such that $\chi=1$ on $\mathop{}\!\mathrm{supp}(\theta)$, then $\overline{A_{\lambda}}=\chi(P_{0,\lambda})\overline{A_{\lambda}}$. In fact, if $f\in D(K_{1})$, then, by (4.26),
    $\overline{A_{\lambda}}f=A_{\lambda}f=F(P_{\lambda})K_{1}f+\frac{1}{2}\left(\theta^{\prime}(P_{0,\lambda})-\theta(P_{0,\lambda})\frac{P_{0,\lambda}}{P_{1,\lambda}^{2}}\right)f=\chi(P_{0,\lambda})A_{\lambda}f=\chi(P_{0,\lambda})\overline{A_{\lambda}}f.$ (4.29)
    This identity extends to $f\in D(\overline{A_{\lambda}})$ by approximating $f$ with elements from $D(K_{1})$ in the graph topology of $\overline{A_{\lambda}}$.
  4. (iv)
    We prove $P_{0,\lambda}\in C^{1}(\overline{A_{\lambda}})$. Utilising $[P_{0,\lambda},\mathop{}\!\mathrm{i}K_{1}]=P_{1,\lambda}$ (see Lemma 4.2), it is easy to verify that, for every $f\in D(K_{1})\cap D(P_{0,\lambda})$,
    $|\langle P_{0,\lambda}f,\overline{A_{\lambda}}f\rangle-\langle\overline{A_{\lambda}}f,P_{0,\lambda}f\rangle|=|\langle f,\theta(P_{0,\lambda})f\rangle|\leq\|\theta(P_{0,\lambda})\|\|f\|^{2}.$ (4.30)
    By approximating $f\in D(\overline{A_{\lambda}})$ in the graph topology of $\overline{A_{\lambda}}$ with elements from $D(K_{1})$, it follows from the previous step that (4.30) is valid for $f\in D(\overline{A_{\lambda}})$. Moreover, $\{f\in D(\overline{A_{\lambda}})\mid(P_{0,\lambda}\pm\mathop{}\!\mathrm{i})^{-1}f\in D(\overline{A_{\lambda}})\}$ contains the core $D(K_{1})$. In fact, if $f\in D(K_{1})$, then
    $\lim_{t\to 0}\frac{1}{\mathop{}\!\mathrm{i}t}(\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tK_{1}}-1)(P_{0,\lambda}\pm\mathop{}\!\mathrm{i})^{-1}f=-\mathop{}\!\mathrm{i}P_{1,\lambda}(P_{0,\lambda}\pm\mathop{}\!\mathrm{i})^{-2}f+(P_{0,\lambda}\pm\mathop{}\!\mathrm{i})^{-1}K_{1}f\in\mathcal{H}.$ (4.31)
    We conclude $P_{0,\lambda}\in C^{1}({\overline{A_{\lambda}}})$ by Proposition 3.2.
  5. (v)
    From the previous step, it follows that the commutator $[P_{0,\lambda},\mathop{}\!\mathrm{i}\overline{A_{\lambda}}]=\theta(P_{0,\lambda})$ is a bounded operator. By similar arguments as in step (iv), we compute the higher order commutators:
    $\mathrm{ad}_{-\mathop{}\!\mathrm{i}\overline{A_{\lambda}}}^{k}(P_{0,\lambda})=\theta_{k-1}(P_{0,\lambda}),\ \theta_{k}(x)=[\theta(x)\partial_{x}]^{k}\theta(x).$ (4.32)
    All commutators are bounded operators; thus, $P_{0,\lambda}\in C^{\infty}(\overline{A_{\lambda}})$. ∎
4.2 Dilation-covariant representations
Let $U:\mathbb{R}^{d}\to\mathfrak{B}(\mathcal{H})$ be a strongly continuous unitary representation of the translation group, whose generators obey the spectrum condition (Assumption 4.1). In this subsection, we assume that $U$ is dilation-covariant, that is, a self-adjoint operator $D$ exists such that, for every $\mu\in\{0,\dots,s\}$ and $t,x\in\mathbb{R}$,
$\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}tD}\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}xP_{\mu}}\mathop{}\!\mathrm{e}^{-\mathop{}\!\mathrm{i}tD}=\mathop{}\!\mathrm{e}^{\mathop{}\!\mathrm{i}\mathop{}\!\mathrm{e}^{-t}xP_{\mu}}.$ (4.33)
We analyse the spectrum of the mass operator $M=\sqrt{P_{0}^{2}-|\mathbf{P}|^{2}}$, which is a well-defined self-adjoint operator due to the spectrum condition.
Proof of Proposition 1.3.
In the terminology of the example subsequent to Definition 3.3, $M$ is $D$-homogeneous due to (4.33), that is, $M\in C^{\infty}(D)$ and
$[M,\mathop{}\!\mathrm{i}D]=M.$ (4.34)
Thus, $M$ obeys a Mourre estimate with conjugate operator $D$ on every open and bounded subset of $[a,\infty)$, $a>0$. Because $a>0$ is arbitrary, it follows from Corollary 3.6 that the spectrum of $M$ is purely absolutely continuous in $(0,\infty)$. ∎
Acknowledgements
I thank Wojciech Dybalski for carefully proofreading the draft of this paper and for assistance preparing the publication. The research leading to these results received funding from National Science Centre, Poland, under the grant ‘Sonata Bis’ 2019/34/E/ST1/00053.